# GM3: A General Physical Model for Micro-Mobility Vehicles

### v1

@@ VI CONCLUSION
This paper introduced the Generalized Micro-Mobility Model (GM3), a tire-level formulation built using the brush-tire model that unifies all types of micro-mobility, including common layouts such as bicycles, scooters, carts, uncommon layouts such as hoverboards and three-wheel platforms, as well as novel layouts. This model captures properties like tire slip, load transfer, and rider/vehicle lean which specialized models like Kinematic Bicycle Model (KBM) miss. To facilitate comparison, we built a model-agnostic simulation framework that decouples wheel layout from dynamics, standardizes control mappings, and supports interactive and scripted simulations. On the Stanford Drone Dataset, GM3 consistently reduced average displacement errors across bicycle, skateboard, and cart modes, while achieving path-level similarity on par with the KBM as measured by the discrete Frechet distance. These observations highlight that grounding dynamics at the tire level results in more realistic local trajectories without sacrificing global path realism. Future work will extend the GM3 to be differentiable for learning, control, and parameter identification, collect instrumented data emphasizing lean and load transfer, study how GM3 enhances interactions between road users, and how it can be used to generate realistic data for autonomous vehicle training and personalized design of autonomous MMVs.

### v2

@@ VI CONCLUSION
This paper introduced the Generalized Micro-Mobility Model (GM3), a tire-level formulation built using the brush-tire model that unifies all types of micro-mobility, including common layouts such as bicycles, scooters, carts, uncommon layouts such as hoverboards and three-wheel platforms, as well as novel layouts. This model captures properties like tire slip, load transfer, and rider/vehicle lean which specialized models like Kinematic Bicycle Model (KBM) miss. Our current implementation supports explicit manual control, lean-to-steer with truck geometry, and differential drive control interfaces, but can be extended to other control methods due to GM3’s modularity. To facilitate comparison, we built a model-agnostic simulation framework that decouples wheel layout from dynamics, standardizes control mappings, and supports interactive and scripted simulations. On the Stanford Drone Dataset, GM3 consistently reduced average displacement errors across bicycle, skateboard, and cart modes, while achieving path-level similarity on par with the KBM as measured by the discrete Fréchet distance. These observations highlight that grounding dynamics at the tire level results in more realistic local trajectories without sacrificing global path realism. Future work will extend the GM3 to be differentiable for learning, control, and parameter identification, collect instrumented data emphasizing lean and load transfer, study how GM3 enhances interactions between road users, and how it can be used to generate realistic data for autonomous vehicle training and personalized design of autonomous MMVs.
ACKNOWLEDGEMENTS
This work is supported in part by Dr. Barry Mersky and Capital One E-Nnovte Endowed Professorships, University of Maryland Distinguished University Professorship, Maryland Transportation Institute Fellowship, National Science Foundation, and ARL-UMD ArtIAMAS Cooperative Agreement.